Solve the following systems of equations using any method.
(b) 2x-1/3y=1/3
3(x+1)-2(y-3)=11

Answer by mananth(16946)

2x-1/3y=1/3
3(2x-1)=3y
6x-3y=0....... 1
..
3(x+1)-2(y-3)=11
3x+3-2y+6=11
3x-2y=2.........2
multiply equation 2 by 2 subtract equation 1
6x-4y=4
..
6x-4y-6x+3y=4
-y=4
y=-4
..
plug the value of y in equation 1
6x+12=0
6x=-12
x=-2
CHECK
6x-3y=0
(6*-2) - (3*-4)=0
Whenever you write fractions please put them in brackets. It makes our life easy